OpenStudy (anonymous):

simplify.. x^2-2x-15 over x+3 divided by x^2+x-30 over x-6

OpenStudy (anonymous):

this is the question ? ((x^2-2x-15)/(x+3))/((x^2+x-30)/(x-6))

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

if so : x^2 - 2x -15 = x^2 -5x +3x -15 = x(x-5) + 3(x-5) = (x-5)(x+3) x^2 + x -30 = x^2 + 6x -5x - 30 = x(x+6) -5(x+6) = (x-5)(x+6) put all together : ((x-5)(x+3)/(x+3))/((x-5)(x+6)/(x-6)) = (x-5)/((x-5)(x+6)/(x-6)) = (x-6)/(x+6)

OpenStudy (anonymous):

thankyou! :)

OpenStudy (anonymous):

you are welcome :)

OpenStudy (lalaly):

\[\huge{\frac{\frac{\cancel{(x+3)}(x-5)}{\cancel{x+3}}}{\frac{(x-5)(x+6)}{x-6}}}\]\[\huge{\frac{\cancel{x-5}}{\frac{\cancel{x-5}(x+6)}{x-6}}}\]\[=\huge{\frac{x-6}{x+6}}\]